You are given the following information about a closed economy with no government: Consumption = 115 + 0.6Y Investment = 550 Use the above information to answer the questions that follow:
Q.4.1 Calculate the value of autonomous spending
1
Expert's answer
2020-09-20T18:41:08-0400

Solution:

Value of autonomous spending = Y

"Y = C+I"

"Y = 115+0.6Y+550"

"Y = 115+550+0.6Y"

"Y = 665+0.6Y"

"Y-0.6Y=665"

"0.4Y=665"

"Y = \frac{665}{0.4}"


"Y = 1663"